1-Kilowatt power was transmitted through a transmission line at 1100 Volts. If transmission line has a resistance of 5 ohm, then power loss due to heat will be approximately:



1
Expert's answer
2021-08-03T11:40:52-0400

The current in the transmission line

"I=\frac{P}{V}=\frac{1000\:\rm W}{1100\:\rm V}=0.909\:\rm A"

The power loss due to heat in the transmission line

"P'=I^2R=\rm (0.909\: A)^2*5\:\Omega=4.13\: W"